**Again-Rules for rounding:

  • If you are calculating in tablets, always round to the nearest 1/2 tablet.
  • When you are calculating IV' rates, round to the nearest whole number.
  • Perform rounding only the end of the calculation.

**Conversions to use:

1 grain =60 mg

15 grain=1 gram

1 pint= 2 cups=480mL

1 quart= 2 4 cups=960mL

1 oz= 30 mL

1 cup = 8 oz = 240 mL

1 quart = 2 pt = 4cups= 960 mL

1cc= 1mL

1tsp= 5mL

1 Tbs= 15 mL

1 gallon= 4 qt = 8 pt

Note:Some drug conversions will use cc(cubic centimeters) This is the same as a mL. 1cc=1mL

QUESTIONS:

1. On the first day after surgery, the patient is started on a clear liquid diet. For breakfast, the patient took 3 ounces of broth, 90 mL of tea, and 2 Tbs of water. For lunch, the patient took 1/4 cup of broth. What is the patients intake for the 0700-1500 shift?______mL

2. The patient is on Intake & Output. For breakfast, the patient took 8 ounces of coffee and 1/2 pint of milk. At lunch, the patient drank 1/2 cup(measured) of juice and 2 T of broth. Calculate the patient’s 8 hr intake for the 0700 – 1500 shift. ______mL.

3. You are asked to make a large quantity of 10% bleach solution. If you need 20mL of total solution, how much bleach do you need? ____mL
Answer #1

1. 3 ounce broth=90 ml

tea=90 ml

2 tbs water= 30 ml

1/4 th cup broth = 60 ml

Total intake=90+90+30+60=270ml

2. 8 ounce coffee= 240ml

Half pint milk= 240 ml

Half cup juice= 120 ml

2 Tbs broth= 30 ml

total intake in shift = 630 ml

3. 2 ml of beach is needed to make 20 ml solution of bleach which is 10%